OpenStudy (jordanloveangel):

5. (3 pt) In which quadrant does the point (–10, 9) lie? A. Quadrant I B. Quadrant II C. Quadrant III D. Quadrant IV

OpenStudy (jordanloveangel):

6. (3 pt) Which answer shows this equation written in standard form? 9 – 4(x – y) = –6x + 6 A. 2x – 4y = 3 B. 2x + 4y = 3 C. –10x + 4y = –3 D. 2x + 4y = –3

OpenStudy (jordanloveangel):

7. (3 pt) Which set of points lies on the given graph? VHS_ALG_S1_07_L203_L303_LQ1-graphic.gif A. (4, 11), (–2, 7), (1, –1) B. (4, 11), (2, –7), (1, 1) C. (–4, 11), (–2, 7), (1, –1) D. (–4, 11), (–2, 7), (1, 1) http://static.k12.com/calms_media/media/213000_213500/213121/1/26e67be17cdf1f48a07be2ef4fbec972478e4772/VHS_ALG_S1_07_L203_L303_LQ1-graphic.gif

OpenStudy (jordanloveangel):

@Ashleyisakitty

OpenStudy (jordanloveangel):

@Catlover5925

OpenStudy (catlover5925):

ok the first one it lies in quad 2, let me solve the second one

OpenStudy (jordanloveangel):

kk

OpenStudy (catlover5925):



OpenStudy (catlover5925):

ok and on the third question for points you have to go over then up to plot them so I started with 4, 11 and it doesn't touch the line at all so we kno A, and B are out so its either C or D, now since 2, -7 are the same on both you need to look at 1, 1 or 1, -1 and your ANSWER is D

OpenStudy (jordanloveangel):

i have like 5 more can u help me and yes it helped @Catlover5925

OpenStudy (catlover5925):

yes I can do you want to open a new strand??

OpenStudy (jordanloveangel):

yes

OpenStudy (catlover5925):

k then tag me